# Concept Complement Bottleneck Model for Interpretable Medical Image Diagnosis

### v1

@@ 3.3.2 Concept Complement Strategy
4 Experiments
4.1 Implementation Details
As for dermoscopic image datasets Derm7pt and Skincon, we follow PCBM [30] and choose the trained Inception-v3 model [40] as the image encoder, and we use pretrained ResNet50 as the image encoder for other two datasets. The frozen pre-trained ClinicalBERT [41] is utilized as the text encoder. In our experiments, the concept adapters and aggregators are set as FCLs. Additionally, we use Adam optimizer for model training. We use the grid method to choose model hyperparameters. For the Derm7pt and BrEaST dataset, the hyperparameters are set as $\lambda_{1}=0.2,\lambda_{2}=10$. For the Skincon dataset, $\lambda_{1}=0.1,\lambda_{2}=5$, and $\lambda_{1}=0.5,\lambda_{2}=10$ for LIDC-IDRI dataset. As for the number of unknown concepts, we can set $n_{u}$ as the number of categories of datasets, since intuitively we need at least the same number of data features as the sample categories to distinguish all categories. To evaluate model performance, we use the Area Under Curve (AUC), ACCuracy (ACC) and F1-score as disease diagnosis evaluation metrics, and the first two metrics are also used to evaluate the concept detection tasks of Derm7pt, Skincon and BrEaST. Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are used for evaluating the concept Regression task of LDIC-IDRI dataset. All of mean and standard deviation results are obtained by five-fold cross-validation in our experiments.
4.2 Comparison Algorithms
4.3 Experimental Results and Analysis
@@
Figure 2: The fine-grained results of the concept detection task on the Derm7pt, BrEaST and Skincon datasets.The results are the means and stds of five-fold cross-validation experiments.
@@
Dataset $n_{u}$ Disease Diagnosis Concept Detection
@@
4.3.1 Disease Diagnosis and Concept Detection
For the Derm7pt dataset, CCBM achieves the best performance in the concept detection task, significantly outperforming the comparison explainable methods with 93.96% AUC, the 88.15% ACC and 85.61% F1-scores. Additionally, CCBM surpasses the black-box model by 1% to 2% on all classification task metrics. Regarding the Skincon dataset, CCBM excels in the concept detection task, particularly with the AUC 20% higher than other competitors, and the best ACC of 91.35%. While maintaining strong concept detection performance, the AUC reaches 84.55%, outperforming other competitors, and the ACC achieves an impressive 80.00%, slightly lower than the best result of 80.24%. The black-box model underperforms on the dataset, with CCBM surpassing it by 5% on AUC, 3% on AUC and 7% on F1-score.
Fig. 2 displays the fine-grained evaluation results of CCBM and two other competitors on the concept detection task. It is evident that CCBM achieves the highest AUC for most concepts across these three datasets. Particularly, our CCBM demonstrates strong performance across all concepts, while CBM exhibits significant misclassification for certain specific concepts compared to its performance on others, such as the BrH and ErY in the Fig. 2(c), the PF and CAL concepts in the Fig. 2(b). This indicates that CCBM treats all concepts more equally and provides more accurate concept detection results than other methods. This advantage stems from the concept adaptation mechanism in our model, which enables the extraction of specific concept features for each concept and the calculation of concept scores in their dedicated channels.
4.3.2 The Impact of the Number of Unknown Concepts on the Model
@@
4.3.3 Abalation Study
The similarity loss $\mathcal{L}_{sim}$ is designed to ensure that the model distinguishes between known and unknown concepts, thereby enhancing performance in disease diagnosis tasks. Essentially, the greater the differentiation between these concepts, the more information the model can assimilate, leading to improved disease diagnosis outcomes. To validate the efficacy of the similarity loss, ablation experiments were performed across the four datasets, with the qualitative results presented in Table 4. It is evident that the similarity loss $\mathcal{L}{sim}$ plays a pivotal role in the model’s performance. Models incorporating the similarity loss $\mathcal{L}{sim}$ outperform those without it across all disease diagnosis metrics, aligning with our expectations. However, models lacking the similarity loss $\mathcal{L}{sim}$ exhibit superior performance in concept detection tasks across the four datasets. The possible reason is, when concepts are less distinct, the model can leverage information from known concepts more effectively, particularly since these concepts are annotated, thus benefiting the concept detection.
4.4 Explanability Analysis
@@
(a) The Results on Derm7pt
(b) The Results on Skincon
(c) The Results on BrEaST
(d) The Results on LIDC-IDRI
(a) The Results on Derm7pt
(b) The Results on Skincon
(c) The Results on BrEaST
(d) The Results on LIDC-IDRI
Figure 5: Visual & Textual Explanations of two images from Derm7pt and LIDC-IDRI, respectively. We visualize the known concepts and unknown concepts (named C1 and C2) of these two examples. The value in brackets represents the truth label of the concept. (For Derm7pt, only the existing concepts are shown. ) Concept scores in red box indicate incorrectly predicted concepts. In the textual explanations, green text denotes correct predictions, while red text highlights incorrect predictions. Blue text presents the insights from the learned unknown concepts.
@@ 4.4.2 Label Efficiency
Obtaining diagnosis labels and annotations for medical images, crucial for improving model interpretability, is often challenging. Hence, ensuring a strong correlation between concepts and model decisions is essential for leveraging labeled data effectively. To validate the label effectiveness of CCBM, we vary the proportions of annotated training data in each five-fold cross-validation experiment while retaining the original full test data for testing. As depicted in Fig. 4, the model performance declines as the Training Data Proportion (TDP) decreases, albeit with only slight drops in the three metrics. Across the Derm7pt, Skincon, and LIDC-IDRI datasets, model performance decreases gradually, with metrics declining by 5% to 10%. The BrEaST dataset exhibits a more pronounced performance decline, attributed to its small training data size. On the LIDC-IDRI dataset, performance degradation remains modest until TDP decreases to 30%. However, a significant performance drop occurs when TDP reaches 10%, deviating from previous trends. These experiments shows CCBM’s efficient utilization of concept and diagnostic labels during training and CCBM demonstrates robust performance when the number of well-annotated samples exceeds approximately 200.
4.4.3 Visual & Textual Explanations for Plausibility
@@ 5 Conclusion
In this paper, we propose a concept complement bottleneck model for interpretable medical image analysis by jointly learning unknown concepts while using known concepts to predict disease. Our model incorporates concept adapters and aggregators with a visual-text concept cross-attention module, creating an fair concept bottleneck model that enhances the precision and effectiveness of disease predictions using known concepts. We also present an effective strategy for learning unknown concepts, aimed at extracting more significant information to enhance model performance. Through comprehensive experiments, we demonstrate that our model achieves superior classification performance in concept detection and disease diagnosis tasks, providing more faithful and understandable explanations. Detailed experiment analysis showcases the effectiveness of the learned unknown concepts. While our study does not delve into the development of new concept textualization, specialization, and generalization, these areas present promising avenues for future exploration.

### v2

@@ 4.3.1 Disease Diagnosis and Concept Detection
For the Derm7pt dataset, CCBM achieves the best performance in the concept detection task, significantly outperforming the comparison explainable methods with 93.96% AUC, 88.15% ACC and 85.61% F1-scores. Additionally, CCBM surpasses the black-box model by 1% to 2% on all classification task metrics. Regarding the Skincon dataset, CCBM excels in the concept detection task, particularly with the AUC 20% higher than other competitors, and the best ACC of 91.35%. While maintaining strong concept detection performance, the AUC reaches 84.55%, outperforming other competitors, and the ACC achieves an impressive 80.00%, slightly lower than the best result of 80.24%. The black-box model underperforms on the dataset, with CCBM surpassing it by 5% on AUC, 3% on AUC and 7% on F1-score.
Fig. 2 displays the fine-grained evaluation results of CCBM and two other competitors on the concept detection task. It is evident that CCBM achieves the highest AUC for most concepts across these three datasets. Particularly, our CCBM demonstrates strong performance across all concepts, while CBM exhibits significant misclassification for certain specific concepts compared to its performance on others, such as the BrH and ErY in the Fig. 2(c), the PF and CAL concepts in the Fig. 2(b). This indicates that CCBM treats all concepts more equally and provides more accurate concept detection results than other methods. Generally, CCBM has higher average AUC and smaller variance for all concepts. This advantage stems from the concept adaptation mechanism in our model, which enables the extraction of specific concept features for each concept and the calculation of concept scores in their dedicated channels.
@@
Figure 2: The fine-grained results of the concept detection task on the Derm7pt, BrEaST and Skincon datasets.The results are the means and stds of five-fold cross-validation experiments. The ”Avg” is the mean and variance of AUC over all concepts.
4.3.2 The Impact of the Number of Unknown Concepts on the Model
@@
Dataset (# of Classes) $n_{u}$ Disease Diagnosis Concept Detection
@@
4.3.3 Ablation Study
@@
The similarity loss $\mathcal{L}_{sim}$ is designed to ensure that the model distinguishes between known and unknown concepts, thereby enhancing performance in disease diagnosis tasks. Essentially, the greater the differentiation between these concepts, the more information the model can assimilate, leading to improved disease diagnosis outcomes. To validate the efficacy of the similarity loss, ablation experiments were performed across the four datasets, with the qualitative results presented in Table 4. It is evident that the similarity loss $\mathcal{L}_{sim}$ plays a pivotal role in the model’s performance. Models incorporating the similarity loss $\mathcal{L}_{sim}$ outperform those without it across all disease diagnosis metrics, aligning with our expectations. However, models lacking the similarity loss $\mathcal{L}_{sim}$ exhibit superior performance in concept detection tasks across the four datasets. The possible reason is, when concepts are less distinct, the model can leverage information from known concepts more effectively, particularly since these concepts are annotated, thus benefiting the concept detection.
4.4 Explanability Analysis
@@ 4.4.2 Label Efficiency
Obtaining diagnosis labels and annotations for medical images, crucial for improving model interpretability, is often challenging. Hence, ensuring a strong correlation between concepts and model decisions is essential for leveraging labeled data effectively. To validate the label effectiveness of CCBM, we vary the proportions of fully annotated training data in each five-fold cross-validation experiment while retaining the original full test data for testing. As depicted in Fig. 4, the model performance declines as the Training Data Proportion (TDP) decreases, albeit with only slight drops in the three metrics. Across the Derm7pt, Skincon, and LIDC-IDRI datasets, model performance decreases gradually, with metrics declining by 5% to 10%. The BrEaST dataset exhibits a more pronounced performance decline, attributed to its small training data size. On the LIDC-IDRI dataset, performance degradation remains modest until TDP decreases to 30%. However, a significant performance drop occurs when TDP reaches 10%, deviating from previous trends. These experiments shows CCBM’s efficient utilization of concept and diagnostic labels during training and CCBM demonstrates robust performance when the number of well-annotated samples exceeds approximately 200.
4.4.3 Visual and Textual Explanations for Plausibility
(a) Derm7pt
(b) Skincon
(c) BrEaST
(d) LIDC-IDRI
Figure 5: Visual and textual explanations of two images from Derm7pt and LIDC-IDRI, respectively. We visualize the known concepts and unknown concepts (named C1 and C2) of these two examples. The value in brackets represents the truth label of the concept. (For Derm7pt, only the existing concepts are shown. ) Concept scores in red box indicate incorrectly predicted concepts. In the textual explanations, green text denotes correct predictions, while red text highlights incorrect predictions. Blue text presents the insights from the learned unknown concepts.
@@ 5 Conclusion
In this paper, we propose a concept complement bottleneck model for interpretable medical image analysis by jointly learning unknown concepts while using known concepts to predict disease. Our model incorporates concept adapters and aggregators with a visual-text concept cross-attention module, creating a fairier concept bottleneck model that enhances the precision and effectiveness of disease predictions using known concepts. We also present an effective strategy for learning unknown concepts, aiming at extracting more significant information to enhance model performance. Through comprehensive experiments, we demonstrate that our model achieves superior classification performance in concept detection and disease diagnosis tasks, providing more faithful and understandable explanations. Detailed experiment analysis showcases the effectiveness of the learned unknown concepts. Our study has not yet explored the textualization, specialization, and generalization of the discovered new concepts, which need further exploration in the future.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data availability Data will be made available on request.
Acknowledgements
This work was supported by the Hong Kong Innovation and Technology Fund (Project No. MHP/002/22), HKUST (Project No. FS111), and the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Reference Number: T45-401/22-N).